If you react 11.45 grams of Cu, how many grams of NO are produced?

 

3Cu + 8HNO→ 3Cu(NO3)2+ 2NO + 4H2O




1
Expert's answer
2021-02-09T03:42:56-0500

11.45 g Cu ×"\frac{1moleCu}{63.546gCu}" = 0.1801 mole Cu


0.1801 mole Cu × "\frac{2moleNO}{3moleCu}" × "\frac{30.01gNO}{1moleNO}"

= 3.60 g NO